OpenStudy (anonymous):

Please help me understand: Simply this expression and state any restrictions on the variables. x+1/x^2+2x-8 - x/4x-8

OpenStudy (anonymous):



OpenStudy (anonymous):

Google...

OpenStudy (anonymous):

have you tried that?

OpenStudy (anonymous):

Yes sir I have,

OpenStudy (anonymous):

I just wanna get an explanation to the question really, I don't just want the answer you know.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

hold on I figure it out

OpenStudy (anonymous):

x = 0.2

OpenStudy (anonymous):

explanation: Assuming you meant (4x - 8) over (x² - 2x), First factor each expressions. You will find some cancelling out. (4x - 8) / (x² - 2x) = 4(x - 2) / x(x - 2) The x - 2 in numerator and denominator cancel, leaving you with 4/x <---ANSWER Looking at the original expression, we don't want denominator to equal 0. x² - 2x = 0 x(x-2) = 0 x = 0, 2 So your expression is undefined at x = 0, 2 <---- ANSWER

OpenStudy (anonymous):

Holy crap, Thank you so much

OpenStudy (anonymous):

no problem...anything else

OpenStudy (anonymous):

No sir, I think that's all.

OpenStudy (anonymous):

alright